# The Cofinality of Generating Families

### v1

@@ 3.2 Upper Bounds - The Covering Number
In personal communication, told to us via Cummings, Gitik thinks ([13]) that the answer to Question A, part 2, for the case of $\mathop{cov}(\kappa)$ is positive, i.e., that the CH is not necessary.
3.3 Lower Bounds - Mod Finite Additivity and Scales
We turn to finding lower bounds of the sampling number, and we start by giving some alternative descriptions of $\mathop{sam}(\kappa)$. Define a general relation $\operatorname{fi}$” by $A\operatorname{fi}B$ if and only if $A$ and $B$ have finite intersection. Note $([\kappa]^{\omega},\operatorname{ii})^{\perp}=([\kappa]^{\omega},\operatorname{fi})$. Define $\kappa]^{\kappa-\omega}=\{\kappa\setminus S:S\in[\kappa]^{\omega}\}$, all co-countably infinite subsets of $\kappa$. For any two sets, $R\subset^{*}S$ if and only if $R\setminus S$ is finite.
Lemma 3.7.
Let $\kappa$ be an uncountable cardinal. The following are Tukey equivalent relations, $([\kappa]^{\kappa-\omega},[\kappa]^{\omega},\not\supset^{*})$, $([\kappa]^{\omega},\operatorname{ii})$, and $([\kappa]^{\omega},[\kappa]^{\kappa-\omega},\not\subset^{*})$. As are their duals, $([\kappa]^{\omega},[\kappa]^{\kappa-\omega},\subset^{*})$, $([\kappa]^{\omega},\operatorname{fi})$, and $([\kappa]^{\kappa-\omega},[\kappa]^{\omega},\supset^{*})$.
Hence, $\mathop{sam}(\kappa)$ is equal to all of the following:
$\begin{matrix}\operatorname{cof}([\kappa]^{\kappa-\omega},[\kappa]^{\omega},\not\supset^{*}),&\operatorname{cof}([\kappa]^{\omega},\operatorname{ii}),&\operatorname{cof}([\kappa]^{\omega},[\kappa]^{\kappa-\omega},\not\subset^{*}),\\ \operatorname{add}([\kappa]^{\omega},[\kappa]^{\kappa-\omega},\subset^{*}),&\operatorname{add}([\kappa]^{\omega},\operatorname{fi}),&\text{ and }\operatorname{add}([\kappa]^{\kappa-\omega},[\kappa]^{\omega},\supset^{*}).\end{matrix}$
Proof.
Observe that complementation in $\kappa$ sets up a bijection between $\kappa]^{\omega}$ and $\kappa]^{\kappa-\omega}$. This bijection establishes the Tukey-equivalence of the first three relations mentioned in the lemma because it is clear from the definitions that, for every $S,T$ from $\kappa]^{\omega}$, we have
[…]
The duals of the first three relations are clearly those stated. To complete the proof recall that the cofinality of a relation equals the additivity of its dual. ∎
Lemma 3.8.
Let $(\kappa_{n})_{n}$ be a sequence of infinite cardinals, each of uncountable cofinality, increasing (not necessarily strictly) to an uncountable cardinal $\kappa$. Then, $(\prod_{n}\kappa_{n},<^{*})\geq_{T}([\kappa]^{\omega},[\kappa]^{{\kappa}-\omega},\subset^{*})$, and hence,
$\mathop{add}(\prod\nolimits_{n}\kappa_{n},<^{*})\leq\mathop{add}([\kappa]^{\omega},[\kappa]^{{\kappa}-\omega},\subset^{*})=\mathop{sam}(\kappa).$
Proof.
The equality in the ‘hence’ part is provided by the preceding lemma. To complete the proof we show $(\prod_{n}\kappa_{n},<^{*})\geq_{T}([\kappa]^{\omega},[\kappa]^{{\kappa}-\omega},\subset^{*})$.
For $f$ in $\prod_{n}\kappa_{n}$, set $\phi_{+}(f)=\kappa\setminus\mathop{ran}(f)$. For $S$ in $\kappa]^{\omega}$, and any $n$, set $\psi_{-}(S)(n)=\sup(S\cap\kappa_{n})$; note that this is a member of $\kappa_{n}$, since $\kappa_{n}$ has uncountable cofinality and $S$ is countable. Now if $\psi_{-}(S)<^{*}f$, then $f(n)\in S\cap\mathop{ran}(f)$ entails that $f(n)\in S\cap\kappa_{n}$, so $f(n)\leq\psi_{-}(S)(n)$, which occurs for only finitely many $n$ - and hence $S\subset^{*}\phi_{+}(f)$. ∎
Note that if a product $\prod_{n\in\omega}\kappa_{n}$ carries a scale of length $\lambda$ regular, then its additivity is $\lambda$. Moving the focus from the sequence to the limit, now suppose that $\kappa$ is a singular cardinal of countable cofinality. Abusing terminology, let us say that $\kappa$ admits a (mod finite) scale of length $\lambda$ if for some strictly increasing sequence $\left<\kappa_{n}:n\in\omega\right$> of regular cardinals cofinal in $\kappa$, the product $\prod_{n\in\omega}\kappa_{n}$ carries a scale of length $\lambda$. By the previous lemma, we conclude:
Corollary 3.9.
Suppose that $\kappa$ is a singular cardinal of countable cofinality. If $\kappa$ admits a scale of regular length $\lambda$ then $\mathop{sam}(\kappa)\geq\lambda$.
The above results give lower bounds on the sampling number. It is not clear to the authors whether taking the supremum of theses lower bounds gives equality. Abusing terminology as above for scales, let us say that an uncountable cardinal $\kappa$, with countable cofinality, admits (mod finite) additivity $\lambda$ if for some increasing sequence $\left<\kappa_{n}:n\in\omega\right$> of regular cardinals cofinal in $\kappa$, the product $\prod_{n\in\omega}\kappa_{n}$ has additivity $\lambda$.
Question B.
  1. 1.
    Is it true in ZFC that $\mathop{sam}(\aleph_{\omega})=\sup\{\lambda:\aleph_{\omega}\text{ admits additivity }\lambda\}$?
  2. 2.
    Is it true in ZFC that $\mathop{sam}(\aleph_{\omega})=\sup\{\lambda:\aleph_{\omega}\text{ admits a scale of length }\lambda\}$?
  3. 3.
    Is it true in ZFC that for every uncountable $\kappa$ of countable cofinality, $\mathop{sam}(\kappa)=\sup\{\lambda:\kappa\text{ admits additivity }\lambda\}$?
  4. 4.
    Is it true in ZFC that for every uncountable $\kappa$ of countable cofinality, $\mathop{sam}(\kappa)=\sup\{\lambda:\kappa\text{ admits a scale of length }\lambda\}$?
[…]
3.4 Jumping Up The Sampling Number
Large $\mathop{sam}(\aleph_{\omega})$ with CH
Many classic models involving Prikry-type forcings contain long mod finite scales. Specifically, Jech has shown in [14] that in Magidor’s classic model [15] – in which, the CH holds and for each $2\leq k<\omega$, we can make $2^{\aleph_{\omega}}=\aleph_{\omega+k}$ – the cardinal $\aleph_{\omega}$ admits a scale of length $2^{\aleph_{\omega}}$. In such a model, we therefore have that $\mathop{sam}(\aleph_{\omega})=\mathop{cov}(\aleph_{\omega})=2^{\aleph_{\omega}}>\aleph_{\omega+1}$.
Theorem 3.10.
If there is a supercompact cardinal then it is consistent that CH and $\mathop{sam}(\aleph_{\omega})=2^{\aleph_{\omega}}>\aleph_{\omega+1}$.
In addition to models where $\mathop{sam}(\aleph_{\omega})=\mathop{cov}(\aleph_{\omega})>\aleph_{\omega+1}$, we also have models where $\mathop{sam}(\aleph_{\omega})=\mathop{cov}(\aleph_{\omega})=\aleph_{\omega+1}$, say if the GCH holds. In both of these cases, however, the two are equal. It is thus a live question of whether we can separate the two and, if so, whether we can make them both above $\aleph_{\omega+1}$.
Question C.
  1. 1.
    Is it true in ZFC that $\mathop{cov}(\aleph_{\omega})=\mathop{sam}(\aleph_{\omega})$?
  2. 2.
    Is it consistent, modulo large cardinals, that $\mathop{sam}(\aleph_{\omega})=\aleph_{\omega+1}<\mathop{cov}(\aleph_{\omega})$?
  3. 3.
    Is it consistent, modulo large cardinals, that $\aleph_{\omega+1}<\mathop{sam}(\aleph_{\omega})<\mathop{cov}(\aleph_{\omega})$?
  4. 4.
    Is it true in ZFC that $\mathop{cov}(\kappa)=\mathop{sam}(\kappa)$, for every uncountable $\kappa$ with countable cofinality?
We know from Corollary 2.4 that for $\kappa$ the first fixed point of the $\aleph$-function the jump from $\kappa$ to $\mathop{cov}(\kappa)$ can be arbitrarily large. Is the same true of the sampling number?
Question D.
Let $\kappa$ be the first fixed point of the $\aleph$-function.
  1. 1.
    Is it true in ZFC that $\mathop{cov}(\kappa)=\mathop{sam}(\kappa)$?
  2. 2.
    Is it consistent, modulo large cardinals, that $\mathop{sam}(\kappa)$ can be arbitrarily large?
@@ 3.4 Jumping Up The Sampling Number
The model above where $\mathop{sam}(\kappa)>\kappa^{+}$ is not directly relevant for our purposes (computing $\mathop{seq}(M)$ and $k(M)$, for $M$ separable metrizable) because CH holds, whereas we need the continuum large while also controlling $\mathfrak{b}$. To correct this we modify via forcing, intending to keep $\mathfrak{b}$ and the relevant sampling numbers unchanged, while boosting the continuum. This leads us to examine the effects of simple forcing notions on the covering and sampling numbers.
We use the following fact about proper, and hence c.c.c., forcings: proper forcing extensions satisfy the countable covering property with respect to $\mathbb{V}$, namely, if $\mathbb{P}$ is proper, then any countable set of ordinals in $\mathbb{V}^{\mathbb{P}}$ is contained in a countable set of ordinals in $\mathbb{V}$. Woodin calls a poset $\mathbb{P}$ weakly proper if and only if it has the countable covering property. We can re-phrase the countable covering property via Tukey quotients.
Lemma 3.11.
If $\mathbb{P}$ is weakly proper and $\kappa$ an uncountable cardinal in $\mathbb{V}$ then in $\mathbb{V}^{\mathbb{P}}$ the identity function witnesses $([\kappa]^{\omega}\cap\mathbb{V},\subseteq)\geq_{T}([\kappa]^{\omega},\subseteq)$.
The above implies that if $\mathbb{P}$ is weakly proper, then the covering number in $V^{\mathbb{P}}$ is no larger than the covering number in $V$. However, it can get smaller.
Example 3.12.
There is a model $V$ and a weakly proper (in fact, $\omega_{1}$-closed) forcing $\mathbb{P}$ that forces the value of $\mathop{cov}(\aleph_{\omega})$ in $V$ to be larger than the value of $\mathop{cov}(\aleph_{\omega})$ in the extension.
To see this, take $V$ to be a model where $\mathop{cov}(\aleph_{\omega})>\aleph_{\omega+1}$ (as above), and let $\mathbb{P}$ be an $\omega_{1}$-closed poset which collapses $2^{\aleph_{\omega}}$ to have size $\aleph_{\omega+1}$.
However, as is well known, a c.c.c. forcing will preserve the covering number.
@@
By Lemma 3.11, $\mathbb{P}$, being weakly proper, forces that $\mu\geq\operatorname{cof}(\dot{X},\subseteq)$.
For the other direction, let $\left<\dot{x}_{\alpha}:\alpha<\lambda\right$> be forced to enumerate a cofinal set in $(\dot{X},\subseteq)$. Apply the c.c.c. of $\mathbb{P}$ to find, for each $\alpha<\lambda$, a ground model set $y_{\alpha}\in[\kappa]^{\omega}$ so that $\Vdash_{\mathbb{P}}\dot{x}_{\alpha}\subseteq\check{y}_{\alpha}$. Then $\left\{y_{\alpha}:\alpha<\lambda\right\}$ is a cofinal set in $([\kappa]^{\omega},\subseteq)$. Thus $\mu\leq\lambda$. ∎
The authors do not know if the sampling number is preserved by c.c.c. forcings. However, a c.c.c. forcing will not decrease it.
@@
Let $\kappa$ be an uncountable cardinal, and let $\mathbb{P}$ be c.c.c. Let $\dot{\lambda}$ denote the $\mathbb{P}$-name for $\mathop{sam}(\kappa)$ in the extension, i.e., the cofinality of $(\dot{X},\operatorname{ii})$, where $\dot{X}$ is the $\mathbb{P}$-name for the set of countable subsets of $\kappa$ in the extension. Then $\mathbb{P}$ forces that $\mathop{sam}^{V}(\kappa)\leq\dot{\lambda}$.
@@
Fix an arbitrary condition $p\in\mathbb{P}$; by extending and relabeling if necessary, we may assume that $p$ decides the value of $\dot{\lambda}$, say as $\lambda$. Let $\left<\dot{x}_{\alpha}:\alpha<\lambda\right$> be a sequence forced by $p$ to be cofinal in $(\dot{X},ii)$. Apply the fact that $\mathbb{P}$ is c.c.c. to find a sequence $\left<y_{\alpha}:\alpha<\lambda\right$> of elements in $\kappa]^{\omega}$ so that for each $\alpha$, $p\Vdash_{\mathbb{P}}\dot{x}_{\alpha}\subseteq\check{y}_{\alpha}$.
We claim that $\left\{y_{\alpha}:\alpha<\lambda\right\}$ is cofinal in $([\kappa]^{\omega},\operatorname{ii})$ in $V$. Fix $x\in[\kappa]^{\omega}$. Let $G$ be an arbitrary $V$-generic over $\mathbb{P}$ containing $p$. In $V[G]$, let $\alpha<\lambda$ so that $x\cap x_{\alpha}$ is infinite; here $x_{\alpha}=\dot{x}_{\alpha}[G]$. Then $x_{\alpha}\subseteq y_{\alpha}$, and hence $x\cap y_{\alpha}$ is infinite in $V[G]$. Since these are in the ground model, $x\cap y_{\alpha}$ is infinite in $V$ as well. ∎
Theorem 3.15.
If there is a supercompact cardinal then it is consistent that $\mathfrak{b}=\mathfrak{d}=\aleph_{1}$, $\mathfrak{c}$ anything we like, and $\mathop{sam}(\aleph_{\omega})=\mathop{cov}(\aleph_{\omega})>\aleph_{\omega+1}$.
@@
Take Magidor’s model from [15] and recall, see Theorem 3.10, in this model the CH holds, and $\mathop{sam}(\aleph_{\omega})=\mathop{cov}(\aleph_{\omega})>\aleph_{\omega+1}$.
Now do random real forcing to keep $\mathfrak{b}$ and $\mathfrak{d}$ equal to $\omega_{1}$, push $\mathfrak{c}$ as large as we like, and not change $\mathop{sam}(\aleph_{\omega})=\mathop{cov}(\aleph_{\omega})>\aleph_{\omega+1}$. This last claim follows from Lemmas 3.14 and 3.13. ∎
Theorem 3.16.
If there is a supercompact cardinal then it is consistent that $\mathfrak{b}=\mathfrak{d}=\aleph_{\omega+1}$, $\mathfrak{c}$ anything we like, and $\mathop{sam}(\aleph_{\omega+1})>\mathfrak{b}$.
Proof.
Take Magidor’s model. Note that $\mathop{sam}(\aleph_{\omega+1})=\mathop{sam}(\aleph_{\omega})>\aleph_{\omega+1}$.
Apply Cohen forcing to get $\mathfrak{b},\mathfrak{d}$ and $\mathfrak{c}$ all equal to $\aleph_{\omega+1}$, without changing $\mathop{sam}(\aleph_{\omega})$ or $\mathop{cov}(\aleph_{\omega})$. Now do random real forcing to leave $\mathfrak{b}$ and $\mathfrak{d}$ unchanged, push $\mathfrak{c}$ as large as we like, and still not change $\mathop{sam}(\aleph_{\omega})$ or $\mathop{cov}(\aleph_{\omega})$. The claimed invariance of $\mathop{sam}(\aleph_{\omega})$ and $\mathop{cov}(\aleph_{\omega})$ follows from Lemmas 3.14 and 3.13, and the initial conditions. ∎
@@ 4.1 How Many Sequences Suffice?
Combining the previous theorem with the Tukey characterization of $\mathbf{seq}(M)$, Theorem 4.1 above, we compute $\mathop{seq}(M)$, for uncountable separable metrizable $M$, in terms of the sampling number of the size of $M$ and $\mathfrak{b}$.
@@
    If $M$ is not locally small then $\mathop{seq}(M)=\mathop{sam}(|M|)\cdot\mathfrak{b}$.
@@
    If $M$ is locally small then $\mathop{seq}(M)=\lim_{\mu<|M|}\mathop{sam}(\mu)\cdot\mathfrak{b}$.
@@
  • We can compute $\mathop{seq}(M)$ from $\mathop{sam}(|M|)$ and $\mathfrak{b}$: $\mathop{seq}(M)=\mathop{sam}(|M|)\cdot\mathfrak{b}$, unless $M$ is locally small when $\mathop{seq}(M)=\lim_{\mu<|M|}\mathop{sam}(\mu)\cdot\mathfrak{b}$.
  • All possible values of $\mathop{seq}(M)$ are attained: if $\kappa\leq\mathfrak{c}$ is uncountable, then there is an $M(\kappa)$ with $\mathop{seq}(M(\kappa))=\mathop{sam}(\kappa)\cdot\mathfrak{b}$; if in addition, $\kappa$ has countable cofinality, then there is an $N(\kappa)$ with $\mathop{seq}(N(\kappa))=\lim_{\mu<\kappa}\mathop{sam}(\mu)\cdot\mathfrak{b}$. This follows a fortiori from Theorem 4.2.
@@
To go deeper, observe that two cases arise depending on the relative sizes of $\mathop{sam}(|M|)$ and $\mathfrak{b}$. If $\mathop{sam}(|M|)\leq\mathfrak{b}$ then $\mathop{seq}(M)=\mathfrak{b}$. In particular, if $\mathfrak{b}=\mathfrak{c}$ then $\mathop{seq}(N)=\mathfrak{c}$ for all uncountable separable metrizable $N$.
Now suppose $\mathfrak{b}\leq\mathop{sam}(|M|)$. In this case the following hold.

### v2

@@ 2.2 PCF Theory
In personal communication, told to us via Cummings, Gitik thinks ([14]) that the answer to Question A, part 2, is positive, i.e., that the CH is not necessary.
@@
The models above where $\mathop{cov}(\kappa)>\kappa^{+}$ is not directly relevant for our purposes (computing $\mathop{seq}(M)$ and $k(M)$, for $M$ separable metrizable) because CH holds, whereas we need the continuum large while also controlling $\mathfrak{b}$. To correct this we modify via forcing, intending to keep $\mathfrak{b}$ and the relevant covering numbers unchanged, while boosting the continuum.
As is well known c.c.c. forcings do not change the covering number.
@@
In particular, we can use random real forcing to push the continuum high, while keeping the bounding number low, and fixing the covering number.
Theorem 2.10.
If there is a supercompact cardinal then it is consistent that $\mathfrak{b}=\mathfrak{d}=\aleph_{1}$, $\mathfrak{c}$ as large as we like, and $\mathop{cov}(\aleph_{\omega})>\aleph_{\omega+1}$.
@@
Take Magidor’s model from [15] and recall in this model the CH holds, and $\mathop{cov}(\aleph_{\omega})>\aleph_{\omega+1}$. Now do random real forcing to keep $\mathfrak{b}$ and $\mathfrak{d}$ equal to $\omega_{1}$, push $\mathfrak{c}$ as large as we like, and not change $\mathop{cov}(\aleph_{\omega})>\aleph_{\omega+1}$. This last claim follows from Lemma 2.9. ∎
A similar argument gives the following.
Theorem 2.11.
If there is a supercompact cardinal then it is consistent that $\mathfrak{b}=\mathfrak{d}=\aleph_{\omega+1}$, $\mathfrak{c}$ as large as we like, and $\mathop{cov}(\aleph_{\omega+1})>\mathfrak{b}$.
While from Gitik’s example, Corollary 2.7, we derive this next scenario.
Theorem 2.12.
If there are sufficiently large cardinals then it is consistent that for $\kappa$ the first fixed point of the $\aleph$-function we have $\mathfrak{b}=\mathfrak{d}=\aleph_{1}$, $\mathop{cov}(\kappa)$ as large as we like, and $\mathfrak{c}$ even larger.
Lower Bounds of the Covering Number - Mod Finite Scales
Below, see Section 3.2, we will want to control the covering number not just of a single cardinal but of families. This will be achieved via lower bounds on the covering number provided by mod finite scales.
Call a set $A$ of cardinals suitable for a cardinal $\kappa$ of countable cofinality if $A$ is countable, every member of $A$ is uncountable, regular and $\sup A=\kappa$. Let $^{*}$ be the mod finite order on $\prod A$, so $f<^{*}g$ if and only if for all but finitely many $\mu$ we have $f(\mu)<g(\mu)$.
@@
If $A$ is suitable for $\kappa$ then $([\kappa]^{\omega},\subseteq)\geq_{T}(\prod A,<^{*})$.
Hence, $\mathop{cov}(\kappa)\geq\sup\{\operatorname{cof}\left(\prod A,<^{*}\right):A\text{ suitable for }\kappa\}$.
@@
For $S$ in $\kappa]^{\omega}$ and $\mu$ in $A$, set $\phi_{+}(S)(\mu)=\sup(S\cap\mu)+1$, and note, since $S$ is countable and $\mu$ has uncountable cofinality, that this is in $\mu$; so $\phi_{+}$ is in $\prod A$. For $f$ in $\prod A$ set $\psi_{-}(f)=\mathop{im}(f)$. Now if $\psi_{-}(f)\subseteq S$ then, for all $\mu\in A$, we have $f(\mu)$ in $S$, so $\phi_{+}(f)(\mu)>\sup(S\cap\mu)\geq f(\mu)$; so $f<^{*}\phi_{+}(S)$, as required. ∎
For sufficiently small $\kappa$ the above lower bound on the covering number is attained. Suppose $\kappa=\aleph_{\alpha}$ is such that $1\leq\alpha<\omega_{1}$ and $\alpha$ has cofinality $\omega$. Then $\kappa$ has a maximum suitable set $A_{\infty}$, namely $A_{\infty}=[\aleph_{1},\kappa]\cap\textsc{REG}$. By a standard PCF fact (see [1]), $\mathop{cov}(\kappa)=\mathop{max\ pcf}(\kappa)=\operatorname{cof}(\prod A_{\infty},<)$; this in turn equals $\operatorname{cof}(\prod A_{\infty},<^{*})$. Thus we have shown:
Lemma 2.14.
If $\kappa=\aleph_{\alpha}$ has countable cofinality and $1\leq\alpha<\omega_{1}$ then
$\mathop{cov}(\kappa)=\sup\{\operatorname{cof}\left(\prod A,<^{*}\right):A\text{ suitable for }\kappa\}.$
A (mod finite) scale in $\prod A$ is a sequence $\vec{f}=\left<f_{\alpha}:\alpha<\lambda\right$> of functions in the product which is increasing mod finite ($\alpha<\beta<\lambda$ implies $f_{\alpha}<^{*}f_{\beta}$) and cofinal mod finite. The cardinal $\lambda$ is called the length of the scale. In Tukey terms, there is a scale of length $\lambda$ in $\prod A$ if and only if $\left(\prod A,<^{*}\right)=_{T}\lambda$. If a product $\prod A$ carries a scale of length $\lambda$ regular, then its cofinality is $\lambda$. Abusing terminology, let us say that $\kappa$ admits a (mod finite) scale of length $\lambda$ if for some $A$ suitable for $\kappa$ the product $\prod A$ carries a scale of length $\lambda$. By Lemma 2.13, we conclude:
Corollary 2.15.
Suppose that $\kappa$ is a singular cardinal of countable cofinality. If $\kappa$ admits a scale of regular length $\lambda$ then $\mathop{cov}(\kappa)\geq\lambda$.
2.3 The Sampling Number
Let $\operatorname{ii}$ be the general relation ($\operatorname{ii}$ for “infinite intersection”) defined as follows: $A\operatorname{ii}B$ if and only if $A\cap B$ is infinite. Define the sampling number of a cardinal $\kappa$ to be $\mathop{sam}(\kappa)=\operatorname{cof}([\kappa]^{\omega},\operatorname{ii})$. That is, $\mathop{sam}(\kappa)$ is the least number of countable subsets of $\kappa$ needed to have infinite intersection with every countably-infinite subset of $\kappa$.
Clearly the sampling is closely related to the covering number. The next result, due to Will Brian [3], tells us that in fact they are equal.
Theorem 2.16 (Brian).
For every infinite cardinal $\kappa$ we have
$([\kappa]^{\omega},\operatorname{ii})=_{T}([\kappa]^{\omega},\subseteq),\quad\text{ and hence }\mathop{sam}(\kappa)=\mathop{cov}(\kappa).$
@@ 2.3 The Sampling Number
First note that a cofinal set for $([\kappa]^{\omega},\subseteq)$ is certainly cofinal in $([\kappa]^{\omega},\operatorname{ii})$, so taking $\phi_{+}$ to be the identity shows $([\kappa]^{\omega},\subseteq)\geq_{T}([\kappa]^{\omega},\operatorname{ii})$.
For the converse, note that $\kappa$ and $\kappa]^{<\omega}$ are bijective, so $([\kappa]^{\omega},\operatorname{ii})$ is isomorphic to $([[\kappa]^{<\omega}]^{\omega},\operatorname{ii})$. We show $([[\kappa]^{<\omega}]^{\omega},\operatorname{ii})\geq_{T}([\kappa]^{\omega},\subseteq)$. To see this, first for each $T$ in $\kappa]^{\omega}$, fix an enumeration, $T=\{t_{i}\}_{i\in\omega}$, and then define $\psi_{-}:[\kappa]^{\omega}\to[[\kappa]^{<\omega}]^{\omega}$ by $\psi_{-}(\{t_{i}\}_{i\in\omega})=\{\{t_{i}\}_{i\leq n}\}:n\in\omega\}$, and $\phi_{+}:[[\kappa]^{<\omega}]^{\omega}\to[\kappa]^{\omega}$ by $\phi_{+}(S)=\bigcup S$. Now if $\psi_{-}(\{t_{i}\}_{i\in\omega})\operatorname{ii}S$, then for infinitely many $n$, $\{t_{i}\}_{i\leq n}\in S$, so $t_{0},\ldots,t_{n}\in\bigcup S=\phi_{+}(S)$, and hence, $\{t_{i}\}_{i\in\omega}\subseteq\phi_{+}(S)$, as required. ∎
@@ 3.1 How Many Sequences Suffice?
Combining the previous theorem with the Tukey characterization of $\mathbf{seq}(M)$, Theorem 3.1 above, we compute $\mathop{seq}(M)$, for uncountable separable metrizable $M$, in terms of the /covering number of the size of $M$ and $\mathfrak{b}$.
@@
    If $M$ is not locally small then $\mathop{seq}(M)=\mathop{cov}(|M|)\cdot\mathfrak{b}$.
@@
    If $M$ is locally small then $\mathop{seq}(M)=\lim_{\mu<|M|}\mathop{cov}(\mu)\cdot\mathfrak{b}$.
@@
  • We can compute $\mathop{seq}(M)$ from $\mathop{cov}(|M|)$ and $\mathfrak{b}$: $\mathop{seq}(M)=\mathop{cov}(|M|)\cdot\mathfrak{b}$, unless $M$ is locally small when $\mathop{seq}(M)=\lim_{\mu<|M|}\mathop{cov}(\mu)\cdot\mathfrak{b}$.
  • All possible values of $\mathop{seq}(M)$ are attained: if $\kappa\leq\mathfrak{c}$ is uncountable, then there is an $M(\kappa)$ with $\mathop{seq}(M(\kappa))=\mathop{cov}(\kappa)\cdot\mathfrak{b}$; if in addition, $\kappa$ has countable cofinality, then there is an $N(\kappa)$ with $\mathop{seq}(N(\kappa))=\lim_{\mu<\kappa}\mathop{cov}(\mu)\cdot\mathfrak{b}$. This follows a fortiori from Theorem 3.2.
@@
To go deeper, observe that two cases arise depending on the relative sizes of $\mathop{cov}(|M|)$ and $\mathfrak{b}$. If $\mathop{cov}(|M|)\leq\mathfrak{b}$ then $\mathop{seq}(M)=\mathfrak{b}$. In particular, if $\mathfrak{b}=\mathfrak{c}$ then $\mathop{seq}(N)=\mathfrak{c}$ for all uncountable separable metrizable $N$.
Now suppose $\mathfrak{b}\leq\mathop{cov}(|M|)$. In this case the following hold.
@@
  • It is consistent, assuming the existence of sufficiently large cardinals, that there is a separable metrizable $M$ of size the first fixed point of the $\aleph$-function, with $\mathop{seq}(M)$ as far above $|M|$ as we like.